Suppose a ball is drawn at random box containing three white and three black balls. After a ball is drawn, it is then replaced and another is drawn. What is the probability that the first four balls drawn, exactly two are white.
1
Expert's answer
2020-06-02T18:33:01-0400

This is a binomial distribution with "n=4, \;p=0.5."

"P(X=2)=C_4^20.5^20.5^2=6*0.25*0.25=0.375."